OpenStudy (anonymous):

The slope of the line below is -1/7. Write the point-slope equation of the line using the coordinates of the labeled point. (3,3) A. y + 3= 1/7(x+3) B. y + 3 = -1/7(x+3) C. y - 3 = -1/7(x-3) D. y - 3 = 1/7(x-3)

OpenStudy (solomonzelman):

point slope form is: \(\large\color{brown}{ y-y_1=m(x-x_1) }\) where: ~ the point is: \(\large\color{brown}{ (x_1,y_1) }\) ~ the slop is: \(\large\color{brown}{ m }\)

OpenStudy (solomonzelman):

you have: ~ a point: \(\large\color{teal}{ (3,3) }\) ~ a slope: \(\large\color{teal}{ -\frac{1}{7} }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ so plug in this (above) given information into the: \(\large\color{teal}{ y-y_1=m(x-x_1) }\)

OpenStudy (anonymous):

so 3 + 3 =-1/7(3-3)? Like that?

OpenStudy (anonymous):

It would be C? @SolomonZelman

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

yes, that is very good!

OpenStudy (solomonzelman):

C